OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = x – 3

OpenStudy (anonymous):

Hello, Your equation is in the slope intercept form: \[y = mx + c\] Where: \(m = slope\) \(c = y-intercept\)

OpenStudy (anonymous):

ok, so do you know the answer

OpenStudy (anonymous):

Moviestarplanet

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my name on their is highschoolcutie1

OpenStudy (anonymous):

here are some choices A. slope: –3, y-intercept: 1 B. slope: 1, y-intercept: –3 C. slope: –1, y-intercept: –3 D. slope: 3, y-intercept: –1

OpenStudy (anonymous):

I do not give out answers directly, if you do not understand my explanation, please tell me what exactly confuses you so I can help.

OpenStudy (anonymous):

lol @Algorithmic that not the answer dido its a other game XD add me on moviestarplanet

OpenStudy (anonymous):

@The_Jokers_wife Ok i will find you.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

If there is no coefficient to a variable in a equation, the coefficient is actually the number one. \[y = mx + c\] \[y = x - 3\] So the slope of you equation is \(1\), as the variable \(x\) is alone. Can you find the y-intercept?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Close, but not exactly. The operator (the minus/negative) symbol should be included.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You are welcome.